# Foldable SuperNets: Scalable Merging of Transformers with Different Initializations and Tasks

### v1

@@ 2.2 Merging transformers
When $r$ is the inner rank, $M_{l}^{1}\in\mathbb{R}^{d_{l}\times r}$, $M_{l}^{2}\in\mathbb{R}^{r\times n\cdot d_{l}}$, and $M_{l}^{\text{diag}}\in\mathbb{R}^{d_{l}\times n\cdot d_{l}}$ is a concatenation of a $n$ diagonal matrices, each with $d_{l}$ learnable parameters. A similar structure is proposed for $U_{l}$. This ensures the number of learnable parameters is linear with the layer width $d_{l}$. Additionally, we found that adding $M_{l}^{\text{diag}}$ is crucial for FS-Merge, as it enables initializing the Foldable SuperNet with strong initializations such as “first” (Eq. 6).
@@
Data and Augmentation. This work addresses a realistic and challenging setting, involving a limited subset of unlabeled samples used for merging. For transformer merges, augmentations (zhang2018mixup) are employed to expand this subset, as commonly done in regular training. Appendix G.1 studies the effect of using augmentation on accuracy.
@@ 3 Results
We evaluate our method on MLPs (Section 3.1), Vision Transformers (Section 3.2), and Text Transformers (Appendix F.3), and show it achieves state-of-the-art results.
@@
Baselines. We compared with “Original Models”, representing the average accuracy of the models to be merged; and Ensemble (ganaie2022ensemble), which averages the models outputs and then applies classification heads. Note that these are not valid merging methods as they use the original models directly. For legitimate merging techniques, comparisons were made with weight averaging (“average”) (wortsman2022model); RegMean (jin2023dataless) which applies a closed-form linear regression solution to each layer; and distillation (hinton2015distilling) which trains a single model to mimic the pre-classification layer features. ZipIt (stoica2024zipit) averages highly correlated neurons, used only in the MLP experiments, as it is inapplicable for transformers. “Opt” (imfeld2023transformer), uses optimal transport (knight2008sinkhorn) for aligning transformers, and “SLERP” (spherical linear interpolation) (shoemake1985animating), utilized only on the ViT case, as they introduced for transformers.
@@
Metrics. As in ZipIt (stoica2024zipit), we evaluate multi-task merged models using two metrics: per-task accuracy and joint accuracy. In per-task accuracy, we calculate the accuracy for each task individually using only the relevant classification head. Then the mean accuracy across all tasks is reported. In joint accuracy, we calculate the accuracy for each task by making predictions based on the maximum score across all classification heads and reporting the mean accuracy across the tasks.
3.1 Merging Multi-Layer Perceptrons
Merge Method Number of Hidden Layers
1 2 3 4 6
Original Models 96.83 $\pm$ 0.13 96.51 $\pm$ 0.19 96.46 $\pm$ 0.23 95.8 $\pm$ 0.4 96.8 $\pm$ 0.5
Ensemble 94.70 $\pm$ 0.95 95.14 $\pm$ 1.12 95.73 $\pm$ 0.22 95.5 $\pm$ 0.1 95.7 $\pm$ 0.7
Average 94.36 $\pm$ 0.76 85.90 $\pm$ 4.46 78.78 $\pm$ 6.72 61.7 $\pm$ 6.3 25.1 $\pm$ 2.7
RegMean 95.90 $\pm$ 0.37 92.97 $\pm$ 2.71 92.11 $\pm$ 1.90 87.7 $\pm$ 3.7 81.5 $\pm$ 2.4
ZipIt 96.35 $\pm$ 0.17 95.75 $\pm$ 0.58 95.43 $\pm$ 0.50 94.5 $\pm$ 0.5 94.0 $\pm$ 2.2
Distillation 93.35 $\pm$ 1.07 93.13 $\pm$ 1.74 93.71 $\pm$ 0.39 93.3 $\pm$ 0.6 90.7 $\pm$ 1.8
FS-M 95.89 $\pm$ 0.03 95.68 $\pm$ 0.19 95.37 $\pm$ 0.33 94.9 $\pm$ 0.4 94.8 $\pm$ 0.8
FS-M, ZipIt init 96.62 $\pm$ 0.08 96.29 $\pm$ 0.24 96.18 $\pm$ 0.20 95.6 $\pm$ 0.5 96.2 $\pm$ 0.8
@@ 3.1 Merging Multi-Layer Perceptrons
Setting. Our evaluation started with a straightforward experiment merging pairs of MLPs. The MLPs were trained on divided MNIST (lecun1998mnist), meaning it was split into two subsets: images with labels 0-4 and images with labels 5-9. These subsets were further divided into training, validation (10% of the training set), and test sets. Two MLPs were trained separately on these subsets, varying in the number of layers and widths, each initialized with a different seed. The goal is to merge these pairs of models. Note that we did not merge the last linear layer, which acts as a classification head.
@@
FS-Merge. Our method was tested in two variants: FS-Merge in the local version (Eq. 3), which trains a Foldable SuperNet for each layer independently; and FS-Merge ZipIt, which initializes the Foldable SuperNet’s $M$ and $U$ as the solutions of ZipIt (stoica2024zipit), and then optimize them using the local version. Data. All merging methods, excluding “Average”, use features from the original models. Thus, we sampled 64 images from each dataset’s training set to generate the necessary features.
@@
Our results indicate that merging deeper models is more challenging, consistent with previous studies (jordan2023repair). Employing the ZipIt initialization, our method establishes a new SOTA for both per-task and joint accuracy, outperforming ensemble in many cases, and nearly matches the accuracy of “Original Models”. For results on more tasks and FS-Merge versions, see Appendix F.1
Table 3: Merging pairs of ViT-B-16 using 16 original images from each training set and 800 augmented images from each dataset. The per-task and joint accuracy on the test set are reported.
@@
Per-task Joint Per-task Joint Per-task Joint #Parameters Optimized
@@
Average 11.48 1.15 3.71 0.84 6.23 3.68 0
SLERP 7.99 1.27 4.10 0.65 6.67 3.24 0
RegMean 8.40 1.69 6.17 1.27 6.44 0.74 0
Opt 4.51 0.75 4.52 0.88 7.01 2.07 0
Distillation 57.31 52.61 62.93 48.12 66.18 63.16  [^1] M
FS-M 63.18 59.28 66.02 49.42 73.43 69.11 13M
3.2 Merging Vision Transformers
Next, we evaluated our method on merging Vision Transformers (ViT), which were initialized differently and trained on distinct tasks— a much more challenging setting. Models and Data. We pre-trained several ViT-B-16 and ViT-L-14 models (dosovitskiy2021an; touvron2021training) on ImageNet-1K (deng2009imagenet). These models were initialized from distinct random seeds and exposed to training data in varying sequences. Then, each differently pre-trained ViT was fine-tuned on downstream tasks. Following (ilharco2023editing), we fine-tuned on Cars (krause20133d), DTD (cimpoi2014describing), EuroSAT (helber2019eurosat), GTSRB (stallkamp2011german), MNIST (lecun1998mnist), RESISC45 (cheng2017remote), SVHN (netzer2011reading), CIFAR10, and CIFAR100 (krizhevsky2009learning). For extended details regarding the pre-training, datasets and finetuning please refer to Appendix D.
In these experiments, the KD baseline also used the “first” initialization, meaning that the student model was initialized from the first model. We found it to outperform any other initialization, including using traditional merging methods such as “RegMean” for initialization and then applying distillation. See Appendix G.2 for full details.
FS-Merge. We use the global version of FS-Merge, when the whole Foldable SuperNet is trained to reconstruct the features of the original models from the layer preceding the classification head. The Foldable SuperNet’s $M$ and $U$ were parametrized as a concatenation of diagonal matrices plus low-rank matrices (Eq. 12). As in the distillation method, the Foldable SuperNet was initialized using the “first” initialization (Eq. 6), as we found it has the best performance (Appendix G.1). FS-M Seq. (Appendix B.4) is a memory and compute-efficient version of FS-Merge, specifically designed for merging a large number of models.
Table 4: Merging groups of 4 ViT-B-16 with 100 original images from the training set and 1000 augmented images from each dataset (a total of 400 original images and 4,000 augmented images). We report the per-task and joint accuracy on the test set. We will denote: C = Cars, D = DTD, E = EuroSAT, G = GTSRB, M = MNIST, R = RESISC45, S = SVHN, C10 = CIFAR10, C100 = CIFAR100.
@@ 3.2 Merging Vision Transformers
Per-task Joint Per-task Joint Per-task Joint #Parameters Optimized
@@
Average 5.40 1.04 3.66 1.35 4.55 0.38 0
SLERP 5.55 1.04 4.88 0.87 7.27 0.86 0
RegMean 6.38 0.61 4.78 0.61 5.60 0.52 0
Opt 5.79 0.24 3.70 0.38 5.75 2.58 0
Distillation 82.09 67.60 67.31 57.67 37.83 31.71  [^1] M
FS-M 84.34 71.17 67.43 55.35 79.48 71.24 60M
FS-M seq. 83.11 70.34 67.55 56.90 73.13 66.03 18M
@@
In Table 3, pairs of ViT-B-16 models fine-tuned on different tasks were merged, in a low-data scenario of using only 16 original images per dataset. An additional 800 augmented images per dataset were created. FS-merge was used with a low rank of 12. To examine the effect of merging a larger number of models, Table 4 shows the results for merging groups of four ViT-B-16 models. FS-merge was used with a low rank of 32, and FS-Merge seq. was used with a low rank of 16. To evaluate the impact of merging larger models, Table 5 presents the results of merging pairs of ViT-L-14 models. FS-merge was used with a low rank of 32. For extended ViT merging results in all these settings and more, and hyperparameter specifics, see Appendix F.2 and Appendix E.2.
Merging Methods DTD, EuroSAT CIFAR100, SVHN Cars, MNIST
Per-task Joint Per-task Joint Per-task Joint #Parameters Optimized
Original models 81.33 - 94.68 - 96.50 - -
Ensemble 77.71 71.54 94.25 77.90 96.23 96.22 -
Average 6.36 1.40 9.18 8.03 5.98 0.08 0
SLERP 5.21 1.58 5.20 2.59 8.48 5.44 0
RegMean 8.66 4.21 5.64 0.47 10.97 0.13 0
Opt 10.33 3.11 4.68 2.45 5.95 5.67 0
Distillation 78.51 75.84 90.91 85.78 91.82 90.58 342M
FS-M 78.60 74.86 91.68 90.92 95.77 95.09 95M
@@
Discussion. As can be seen, FS-Merge outperforming all other merging methods in most cases, and even surpassing ensembles in some cases. This holds for both per-task and joint accuracy across all settings, despite using fewer learnable parameters than distillation. Additionally, it is evident that all local and simple methods (such as Average, SLERP, RegMean, and “Opt”) completely fail to effectively merge ViTs in this challenging setting, resulting in a merged model that performs comparably to a random guess. We also find that FS-Merge achieves SOTA results when merging BERTs on NLP tasks, as can be seen in Appendix F.3.
Our experiments indicate that in the ViT case, initialization is crucial for FS-Merge as it does not converge when initialized randomly (and see Appendix H.2). Specifically, using the “first” initialization in Foldable SuperNet, and KD, not only improves the accuracy of the first task but also enhances accuracy across all tasks. For an ablation study in the matter, see Appendix G.1 and Appendix G.2.
Merging Complexity. FS-Merge and KD are more computationally intensive than standard merging methods such as Averaging and RegMean, which fail catastrophically in our setting. Notably, the accuracy of both FS-Merge and KD cannot be improved simply by longer training (i.e. more resources), as this causes overfitting in our data-scarce regime (this explains why the longest training duration was not identified as optimal in our hyperparameter search). When merging two models, FS-Merge and distillation have comparable resource usage. However, when merging multiple models, the resource usage gap between FS-Merge and distillation becomes more significant. To address this, we propose FS-Merge seq., which is comparable to distillation’s resource use while outperforming it in terms of accuracy. Therefore, FS-Merge is recommended for optimal test accuracy, given sufficient computational resources and limited data. If resources are more limited, FS-Merge seq. should be the method of choice. Refer to Appendix C for the full details.
@@ 3.3 Number of original training images
We examine the impact of varying $|D|$, the number of images taken from the training datasets of the models to be merged (“original images”), used to create features. We varied $|D|$ from 16 to 1024. Augmented images were created to ensure the total number of images per dataset reached 1024, thus maintaining a consistent dataset size. Pairs of ViT-B-16 models were merged using Ensemble, Distillation, and FS-Merge with low rank of 24. The per-task and joint accuracies on the test set are presented in Figure 4. For additional details and experiments, refer to Appendix F.4.
@@
As observed, Distillation underperforms with few original images, while FS-Merge excels. Increasing original images enhances all techniques, reducing the performance gap. With enough data, merging methods can sometimes surpass ensemble performance, which is often considered as a “gold-standard method” in merging and multitask articles.
We report that FS-Merge and Distillation achieve perfect fit on the original images in all cases. We argue that low-rank FS-Merge attains better generalization than distillation due to a useful inductive bias, which constrains the merged model to be a low-rank weighted average of the original model’s weights (as determined by the Foldable SuperNet).
4 Conclusion

### v2

@@ 2.2 Merging transformers
When $r$ is the inner rank, $M_{l}^{1}\in\mathbb{R}^{d_{l}\times r}$, $M_{l}^{2}\in\mathbb{R}^{r\times n\cdot d_{l}}$, and $M_{l}^{\text{diag}}\in\mathbb{R}^{d_{l}\times n\cdot d_{l}}$ is a concatenation of a $n$ diagonal matrices, each with $d_{l}$ learnable parameters. A similar structure is proposed for $U_{l}$. This ensures the number of learnable parameters is linear with the layer width $d_{l}$. We found that adding $M_{l}^{\text{diag}}$ is crucial for FS-Merge, as it enables initializing the Foldable SuperNet with strong initializations such as “first” (Eq. 6).
@@
Data and Augmentation. This work addresses a realistic setting, involving a limited subset of unlabeled samples used for merging. For transformer merges, augmentations (Zhang et al., 2018) are employed to expand this subset, as commonly done in regular training. Appendix G.1 studies the effect of using augmentation on accuracy.
Expressive power. As shown in Section 3.2, FS-Merge significantly surpasses traditional merging methods. We explain this with theoretical results demonstrating its superior expressivity over the baselines.
Theorem 2.1 (Informal).
FS-Merge can exactly reproduce any target MLP model when merging models with sufficient rank conditions. Moreover, FS-Merge strictly generalizes previous merging methods like RegMean and Git Re-Basin, in the context of merging MLPs.
See Appendix H for the formal statement and conditions.
Model Number of Hidden Layers
1 2 3 4 6
Two models 96.83 $\pm$ 0.13 96.51 $\pm$ 0.19 96.46 $\pm$ 0.23 95.86 $\pm$ 0.40 96.81 $\pm$ 0.58
Ensemble 94.70 $\pm$ 0.95 95.14 $\pm$ 1.12 95.73 $\pm$ 0.22 95.58 $\pm$ 0.12 95.78 $\pm$ 0.73
average 94.36 $\pm$ 0.76 85.90 $\pm$ 4.46 78.78 $\pm$ 6.72 61.76 $\pm$ 6.36 25.12 $\pm$ 2.70
RegMean 95.90 $\pm$ 0.37 92.97 $\pm$ 2.71 92.11 $\pm$ 1.90 87.79 $\pm$ 3.77 81.55 $\pm$ 2.49
ZipIt 96.35 $\pm$ 0.17 95.75 $\pm$ 0.58 95.43 $\pm$ 0.50 94.59 $\pm$ 0.50 94.03 $\pm$ 2.27
Distillation 93.35 $\pm$ 1.07 93.13 $\pm$ 1.74 93.71 $\pm$ 0.39 93.38 $\pm$ 0.67 90.79 $\pm$ 1.89
FS-M 95.89 $\pm$ 0.03 95.68 $\pm$ 0.19 95.37 $\pm$ 0.33 94.94 $\pm$ 0.43 94.89 $\pm$ 0.89
FS-M, ZipIt init 96.62 $\pm$ 0.08 96.29 $\pm$ 0.24 96.18 $\pm$ 0.20 95.63 $\pm$ 0.52 96.22 $\pm$ 0.84
@@ 3 Results
We evaluate our method on MLPs (Section 3.1), Vision Transformers (Section 3.2), and Text Transformers (Section 3.3), and show it achieves SOTA results.
@@
Baselines. We compared with “Original Models”, representing the average accuracy of the models to be merged; and Ensemble (Ganaie et al., 2022), which averages the models outputs and then applies classification heads. Note that these are not valid merging methods as they use the original models directly. For legitimate merging techniques, comparisons were made with weight averaging (Wortsman et al., 2022); RegMean (Jin et al., 2023) which applies a closed-form linear regression solution to each layer; and distillation (Hinton et al., 2015) which trains a single model to mimic the pre-classification layer features. ZipIt (Stoica et al., 2024) averages highly correlated neurons, used only in the MLP experiments, as it is inapplicable for transformers. “Opt” (Imfeld et al., 2023), uses optimal transport for aligning transformers, and “SLERP” (Shoemake, 1985), utilized only on the ViT case, as they introduced for transformers.
Data. When merging the models, a small unlabeled subset of their original training data is used. In the case of merging ViTs, this subset is further extended using augmentations. For a fair comparison, all baselines use the same amount of original and augmented data.
@@
Metrics. Following ZipIt (Stoica et al., 2024), we evaluate multi-task merged models using per-task and joint accuracy. Per-task accuracy is calculated by evaluating each task individually with its relevant classification head and reporting the mean accuracy across all tasks. Joint accuracy is calculated by making predictions for each task based on the maximum score across all classification heads and reporting the mean accuracy across tasks.
3.1 Warm-up: Merging Multi-Layer Perceptrons
@@
Setting. MNIST (LeCun, 1998) was split into two subsets: images with labels 0–4 and images with labels 5–9. These subsets were further divided into training, validation, and test sets. Two MLPs were trained separately on these subsets, varying in the number of layers and widths, each initialized with a different seed. The goal was to merge these pairs of models, excluding the last linear layer, which serves as the classification head.
@@
FS-Merge. Our method was tested in two variants: the local version of FS-Merge (Eq. 3), which trains a Foldable SuperNet for each layer independently; and FS-Merge ZipIt, which initializes the Foldable SuperNet’s $M$ and $U$ as the solutions of ZipIt (Stoica et al., 2024), and then optimize them using the local version. Data. All merging methods, excluding “Average”, use features from the original models. Thus, we sampled 64 images from each dataset’s training set to generate the necessary features.
@@
Our results indicate that merging deeper models is more challenging, consistent with previous studies (Jordan et al., 2023). Employing the ZipIt initialization, our method establishes a new SOTA for both per-task and joint accuracy, outperforming ensemble in many cases, and nearly matches the accuracy of “Original Models”. For results on more tasks and FS-Merge versions, see Appendix F.1
3.2 Merging Vision Transformers
Next, we evaluated our method on merging Vision Transformers (ViT) initialized differently and trained on distinct tasks, a much more challenging problem.
Models and Data. Several ViT-B-16 and ViT-L-14 models (Dosovitskiy et al., 2021; Touvron et al., 2021) were pre-trained on ImageNet-1K (Deng et al., 2009), using distinct random seeds. Then, each differently pre-trained ViT was fine-tuned on downstream tasks. Following (Ilharco et al., 2023), we fine-tuned on Cars (Krause et al., 2013), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), GTSRB (Stallkamp et al., 2011), MNIST (LeCun, 1998), RESISC45 (Cheng et al., 2017), SVHN (Netzer et al., 2011), CIFAR10, and CIFAR100 (Krizhevsky et al., 2009). For extended details regarding the setting refer to Appendix D.
In these experiments, the KD baseline aims to reconstruct the features from the last layer of the models to be merged, using the “first” initialization, meaning that the student model was initialized from the first model. We tested additional KD variants, such as a baseline that attempts to mimic the internal features of the models (Appendix G.5), and baselines where the student model is initialized using different merging methods (such as RegMean) before continuing training (Appendix G.2). Nevertheless, we found that our KD baseline, which uses only the last-layer features with the “first” initialization, outperforms all other variants. The KD LoRA is identical to the KD baseline, except that it trains only LoRA Hu et al. (2022) adapters with a low rank $r$ for each linear layer in the model.
Table 3: Merging pairs of ViT-B-16 using 16 original images from each training set and 800 augmented images from each dataset. The per-task and joint accuracy on the test set are reported
@@ 3.2 Merging Vision Transformers
Per-task Joint Per-task Joint Per-task Joint Learnable Parameters
@@
Average 11.48 1.15 3.71 0.84 6.23 3.68
SLERP 7.99 1.27 4.10 0.65 6.67 3.24
RegMean 8.40 1.69 6.17 1.27 6.44 0.74
Opt 4.51 0.75 4.52 0.88 7.01 2.07
Distillation 57.31 52.61 62.93 48.12 66.18 63.16
Distillation, LoRA $r=4$ 57.23 51.83 58.81 49.83 62.17 57.04
Distillation, LoRA $r=12$ 55.07 48.28 59.06 47.59 64.16 58.66
FS-M, $r=12$ 62.64 57.42 66.50 53.98 73.77 69.26
Table 4: Merging groups of 4 ViT-B-16 using 100 original and 1000 augmented images per training dataset (a total of 400 original images and 4,000 augmented images). We will denote: C = Cars, D = DTD, E = EuroSAT, G = GTSRB, M = MNIST, R = RESISC45, S = SVHN, C10 = CIFAR10, C100 = CIFAR100.
@@
Per-task Joint Per-task Joint Per-task Joint Learnable Parameters
@@
Average 5.40 1.04 3.66 1.35 4.55 0.38
SLERP 5.55 1.04 4.88 0.87 7.27 0.86
RegMean 6.38 0.61 4.78 0.61 5.60 0.52
Opt 5.79 0.24 3.70 0.38 5.75 2.58
Distillation 82.09 67.60 67.31 57.67 45.75 40.62
Distillation, LoRA $r=4$ 72.92 54.73 57.91 46.36 75.76 70.34
Distillation, LoRA $r=12$ 77.43 60.18 63.56 52.81 75.52 69.29
Distillation, LoRA $r=24$ 78.58 61.47 63.61 52.93 68.75 64.31
FS-M, $r=24$ 84.58 70.92 68.41 56.40 79.71 73.98
FS-M seq., $r=16$ 83.90 71.07 68.96 58.48 73.57 68.31
Table 5: We merge pairs of ViT-B-16 models using 16 original and 800 augmented images per dataset. Accuracy is reported on the corrupted test sets of CIFAR100-C, with mean and standard deviation computed over the five severity levels.
Corruption type CIFAR100 ViT-B-16 CIFAR100, RESISC45 CIFAR100, EuroSAT
Distillation FS-Merge Distillation FS-Merge
Snow 71.9 $\pm$ 5.1 63.4 $\pm$ 5.5 71.2 $\pm$ 5.2 64.1 $\pm$ 5.2 69.6 $\pm$ 5.6
Frost 69.6 $\pm$ 8.8 61.2 $\pm$ 8.9 69.0 $\pm$ 8.9 62.2 $\pm$ 9.2 64.6 $\pm$ 10.9
Fog 75.7 $\pm$ 11.2 66.4 $\pm$ 13.1 75.2 $\pm$ 11.2 68.5 $\pm$ 12.4 74.4 $\pm$ 11.3
Brightness 82.8 $\pm$ 2.9 75.4 $\pm$ 3.6 82.4 $\pm$ 3.0 76.4 $\pm$ 3.4 81.8 $\pm$ 3.0
Gaussian noise 24.9 $\pm$ 16.8 22.8 $\pm$ 14.3 24.2 $\pm$ 17.0 22.4 $\pm$ 14.8 21.5 $\pm$ 15.8
Shot noise 33.7 $\pm$ 21.9 30.3 $\pm$ 18.6 33.0 $\pm$ 22.0 30.6 $\pm$ 19.2 30.0 $\pm$ 21.1
Impulse noise 30.0 $\pm$ 20.3 21.8 $\pm$ 15.9 28.6 $\pm$ 19.8 20.3 $\pm$ 17.2 26.3 $\pm$ 20.5
Contrast 69.3 $\pm$ 20.7 61.2 $\pm$ 21.1 69.2 $\pm$ 20.5 62.9 $\pm$ 21.1 66.4 $\pm$ 22.3
Elastic transform 70.5 $\pm$ 12.1 60.1 $\pm$ 12.0 69.9 $\pm$ 12.1 62.6 $\pm$ 11.9 69.3 $\pm$ 12.2
Pixelate 62.9 $\pm$ 22.0 58.1 $\pm$ 18.4 61.5 $\pm$ 22.7 58.4 $\pm$ 19.9 61.4 $\pm$ 21.8
Jpeg compression 60.2 $\pm$ 6.7 51.3 $\pm$ 6.2 59.8 $\pm$ 6.7 53.1 $\pm$ 6.2 59.0 $\pm$ 6.8
Defocus blur 76.3 $\pm$ 10.7 66.3 $\pm$ 13.3 75.9 $\pm$ 10.7 69.0 $\pm$ 12.1 75.3 $\pm$ 10.7
Motion blur 69.2 $\pm$ 8.6 57.7 $\pm$ 9.8 68.9 $\pm$ 8.6 60.9 $\pm$ 9.4 67.8 $\pm$ 8.6
Zoom blur 72.9 $\pm$ 6.1 61.5 $\pm$ 7.0 72.6 $\pm$ 6.1 64.5 $\pm$ 7.0 71.9 $\pm$ 6.1
@@
FS-Merge. The global version of FS-Merge was used (Eq. 4), with the Foldable SuperNet’s $M$ and $U$ parametrized as a concatenation of diagonal matrices plus low-rank matrices (Eq. 7). The Foldable SuperNet was initialized using the “first” initialization (Eq. 6), which showed the best performance (as demonstrated in Appendix G.1). FS-Merge Seq. (Appendix B.4) is a memory and compute-efficient variant of FS-Merge, designed for merging a large number of models. In Table 3, pairs of ViT-B-16 models fine-tuned on different tasks were merged in a low-data scenario of using only 16 original images and 800 augmented images per dataset. To study the impact of merging a larger number of models, Table 4 shows results for merging groups of four ViT-B-16 models. For additional results and hyperparameter details, refer to Appendices F.2 and E.2.
@@
As can be seen, FS-Merge outperforms all other merging methods, and even surpasses ensembles in some cases. This holds for both per-task and joint accuracy across all settings. Additionally, it is evident that all local and simple methods (such as Average, SLERP, RegMean, and “Opt”) completely fail to effectively merge ViTs in this challenging setting, resulting in a merged model that performs comparably to a random guess. Similar results were observed when merging larger ViTs (Appendix F.2) and ViTs trained with different pretraining strategies (Appendix F.3).
Robustness evaluation. In the next section, the robustness of merging methods under distribution shifts was evaluated. The ViT-B-16 model fine-tuned on CIFAR100 was merged with other ViT-B-16 models using KD and FS-Merge. The merged models were then tested on CIFAR100-C (Hendrycks & Dietterich, 2019), a dataset that introduces various corruptions at five severity levels to the CIFAR100 test set. For each corruption type, the mean and standard deviation across the severity levels were calculated (Table 5). The performance of the original ViT-B-16 model trained on CIFAR100 is also included as an upper bound for the merged models.
As demonstrated, FS-Merge consistently outperforms KD across corruption types and severity levels, achieving performance close to the original ViT-B-16 model trained on CIFAR100. This highlights its effectiveness in producing signficantly more robust merged models.
@@
Number of original training images. We examine the impact of varying $|D|$, the number of images from the training datasets of the models to be merged (“original images”), which are used to create features. $|D|$ was varied from 16 to 1024. Augmented images were created to ensure the total number of images per dataset reached 1024, maintaining a consistent dataset size. Pairs of ViT-B-16 models were merged using Ensemble, KD and FS-Merge. The per-task and joint accuracies on the test set are presented in Figure 4. For additional details and experiments, see Appendix F.8.
@@
As observed, Distillation underperforms with few original images, while FS-Merge excels. Increasing original images enhances both techniques, reducing the performance gap. With enough data, merging methods can sometimes surpass ensemble performance, which is often considered as a “gold-standard method” in merging and multitask articles.
3.3 Merging Text Transformers
In this series of experiments, we aimed to evaluate FS-Merge on a different modality: merging text transformers. We fine-tuned differently initialized BERT models on distinct classification task from the GLUE dataset (Wang et al., 2019). Then, we applied KD and FS-Merge to merge pairs of them, utilizing 200 data points from each training set to create features. In Table 6, we report the per-task test accuracy of these experiments. We observe that FS-Merge outperforms KD, similar to the ViT case. For details and additional results, see Appendix F.7.
Table 6: Merging pairs of BERTs with 200 data points from the training set of each dataset. We report the test set per-task accuracy.
Tasks Original models Distillation FS-Merge
RTE, QQP 79.49 65.00 68.88
MNLI, MRPC 85.73 76.87 78.29
MRPC, QNLI 89.02 80.67 79.33
SST-2, RTE 80.15 75.45 76.00
MNLI, SST-2 88.39 82.81 83.24
RTE, QNLI 79.75 66.40 69.69
QNLI, QQP 91.01 82.40 84.75
QQP, RTE 79.49 73.30 75.62
SST-2, QNLI 91.68 82.92 84.39
MRPC, RTE 77.49 68.70 68.37
QNLI, MNLI 87.99 71.83 73.79
QQP, SST-2 91.41 87.49 87.94
MNLI, QQP 87.7 78.74 79.22
RTE, MRPC 77.49 63.46 66.05
4 Discussion
Merging Complexity. FS-Merge and KD are more computationally intensive than standard merging methods such as Averaging and RegMean, which fail catastrophically in our setting. When merging two models, FS-Merge and KD have comparable resource usage. However, when merging multiple models, the resource usage gap between FS-Merge and KD becomes more significant. To address this, we propose FS-Merge seq., which is comparable to KD’s resource use while outperforming it in terms of accuracy. Therefore, FS-Merge is recommended for optimal test accuracy, given sufficient computational resources and limited data. If resources are more limited, FS-Merge seq. should be the method of choice. Refer to Appendix C for the full details.
Why does FS-Merge work well? We attribute FS-Merge’s superior performance to three key aspects. First, FS-Merge and KD offer greater expressiveness than traditional methods (Appendix H). Second, in contrast to the traditional methods, FS-Merge and KD use a global objective, which our experiments show is crucial in this challenging setting (see Appendix G.3). Lastly, we believe FS-Merge generalizes better than KD because it ha s an useful inductive bias that constrains the merged model to be a low-rank weighted average of the original models’ weights (as determined by the Foldable SuperNet).